Let 𝑋1, … , 𝑋𝑛 be iid Gamma(1, 1/𝛼) RVs

(a) Show that the estimator T(𝑋1, … , 𝑋𝑛) = (n-1)/n𝑋̅ is the UMVUE for 𝛼 withΒ 

variance 𝛼2/(𝑛 βˆ’ 2)

(b)Show that the minimum variance from CR inequality is 𝛼2/𝑛


1
Expert's answer
2021-12-17T10:40:49-0500

Solution:

(a):

"X_i\sim G(1,1/\alpha)" are iid for all i=1,..,n

"f_{x_i}(x_i)=\alpha^{-x_i \alpha} ; x_i>0
\\ f( \underline{X})=\Pi_{i=1}^nf_{x_i}(x_i)
\\=\alpha^ne^{-\Sigma x_i \alpha_i}
\\=A(\alpha)h(\underline x)e^{\eta(\alpha)T(\underline x)}
\\A(\alpha)=\alpha^n, h(\underline x)=1,\eta(\alpha)=-\alpha,T'(\underline x)=\Sigma x_i"

"\therefore f( \underline{X})" is from exponential family.

Now, "\Sigma x_i\sim G(n,1/\alpha)"

"\therefore E(\dfrac1{\Sigma x_i})=\int_0^{\infty}\dfrac 1t. \dfrac{\alpha^nt^{n-1}e^{-t \alpha}}{\Gamma(n)}dt
\\=\dfrac{\Gamma(n-1)}{\Gamma(n)}\alpha
\\=\dfrac{\alpha}{n-1}
\\\therefore E(\dfrac{n-1}{n\bar X})=\alpha"

So, "T(\underline X)=\dfrac{n-1}{n\bar X}" is unbiased for "\alpha".

"\therefore E(T(\underline X)|T'(\underline X)=\Sigma x_i)" is UMVUE from Lehmann Scheffe theorem.

"\therefore T(\underline X)=\dfrac{n-1}{n\bar X}" is UMVUE for "\alpha".

"\therefore E(\dfrac1{(\Sigma x_i)^2})=\int_0^{\infty}\dfrac 1{t^2}. \dfrac{\alpha^nt^{n-1}e^{-t \alpha}}{\Gamma(n)}dt
\\=\dfrac{\Gamma(n-2)}{\Gamma(n)}\alpha^2
\\=\dfrac{\alpha^2}{(n-1)(n-2)}"

"V(\dfrac{n-1}{n\bar X})=(n-1)^2V(\dfrac1{(\Sigma x_i)^2})
\\=(n-1)^2[E(\dfrac1{(\Sigma x_i)^2})-E^2(\dfrac1{\Sigma x_i})]
\\=(n-1)^2[\dfrac{\alpha^2}{(n-1)(n-2)}
-\dfrac{\alpha^2}{(n-1)^2}
]
\\=\dfrac{\alpha^2}{n-1}"

(b):

"f_X(x)=\alpha e^{-\alpha x};x>0\ \& \ g(\theta)=\alpha \Rightarrow g'(\theta)=1
\\L(f_X(x))=\alpha^n e^{-\alpha \Sigma x_i}
\\ \log L(f_X(x))=n \log \alpha-\alpha \Sigma x_i
\\ \dfrac{\partial }{\partial \alpha}\log L(f_X(x))=\dfrac{n}{\alpha}-\Sigma x_i
\\ \dfrac{\partial^2 }{\partial \alpha^2}\log L(f_X(x))=\dfrac{-n}{\alpha^2}
\\ I(\theta)=-E[\dfrac{\partial^2 }{\partial \alpha^2}\log L(f_X(x))]
\\=-E[\dfrac{-n}{\alpha^2}]=\dfrac{n}{\alpha^2}"

And CR inequality "=CR=\dfrac{[g'(\theta)]^2}{I(\theta)}"

Here "g'(\theta)=1"

So, "CR=\dfrac{1}{\dfrac n{\alpha^2}}=\dfrac{\alpha^2}{n}"

Hence, proved.